OpenStudy (anonymous):

Please check my answer :DD I got C! :D @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you used A = P*(1+r)^t I bet huh?

jimthompson5910 (jim_thompson5910):

if so then you used the wrong formula notice how it says "continuously"

jimthompson5910 (jim_thompson5910):

A = P*(1+r)^t only applies to annual growth or decay

OpenStudy (anonymous):

:( Here we go again ! >.< I used i= prt and that didn't work either!

jimthompson5910 (jim_thompson5910):

that's only for simple interest

jimthompson5910 (jim_thompson5910):

You'd use A = P*e^(r*t)

OpenStudy (anonymous):

e?

jimthompson5910 (jim_thompson5910):

'e' is a constant like pi e = 2.71828182846 and that expansion goes on forever without a known pattern

OpenStudy (anonymous):

Oh, that e- yes!! lol I thought you meant a variable..

jimthompson5910 (jim_thompson5910):

e is rarely (if ever) used as a variable

OpenStudy (anonymous):

There's always that 1 time, you never know! 71427.46

jimthompson5910 (jim_thompson5910):

that's true

jimthompson5910 (jim_thompson5910):

71,472.46 is correct

OpenStudy (anonymous):

That's all the steps or did I miss something ? :/ If not, that's the final answer then? :)

jimthompson5910 (jim_thompson5910):

yes that's the answer

OpenStudy (anonymous):

Thank you! :)

jimthompson5910 (jim_thompson5910):

np